OpenStudy (anonymous):

hello , please i need help on Q2 , Q5 and the formula of Problem 5 in Q1 : Vth ?!!!

OpenStudy (anonymous):

Q1P5: Vth=VS*RON/(RON+RL)

OpenStudy (anonymous):

What is RL?

OpenStudy (anonymous):

sorry it is not RL but R

OpenStudy (anonymous):

Q5P1:First calculate i through the inductor by replacing the inductor by a short..now il(t)=i*(1-e^(-t/tau)) tau=L/((R1+R2)||R3) t=given value Q5P2:calculate v across cap with c as open..now vc(t)=v*e^(-t/tau)..tau=((R1+R2)||R3)*C FOR BOTH t<0 and t=0+ subs t=0 in tat equt...

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

what is R value

OpenStudy (anonymous):

tell me pzl...

OpenStudy (anonymous):

@kartkn please how to calculate i in Q5 , and v in the 2nd part of Q5

OpenStudy (anonymous):

L=1H, R1=12Ω, R2=9Ω, and R3=6Ω for t=300.0ms Part 2 C=300.0mF, R1=12Ω, R2=9Ω, and R3=6Ω. please some help